### v1

# DR-MPC: Deep Residual Model Predictive Control for Real-world Social Navigation

Abstract
How can a robot safely navigate around people exhibiting complex motion patterns? Reinforcement Learning (RL) or Deep RL (DRL) in simulation holds some promise, although much prior work relies on simulators that fail to precisely capture the nuances of real human motion. To address this gap, we propose Deep Residual Model Predictive Control (DR-MPC), a method to enable robots to quickly and safely perform DRL from real-world crowd navigation data. By blending MPC with model-free DRL, DR-MPC overcomes the traditional DRL challenges of large data requirements and unsafe initial behavior. DR-MPC is initialized with MPC-based path tracking, and gradually learns to interact more effectively with humans. To further accelerate learning, a safety component estimates when the robot encounters out-of-distribution states and guides it away from likely collisions. In simulation, we show that DR-MPC substantially outperforms prior work, including traditional DRL and residual DRL models. Real-world experiments show our approach successfully enables a robot to navigate a variety of crowded situations with few errors using less than 4 hours of training data (video: https://youtu.be/GUZlGBk60uY, code: upon acceptance).
Index Terms:
Social HRI, Reinforcement Learning, Model Predictive Control, Real-world Robotics, Motion Planning.
I Introduction
Achieving reliable robotic navigation remains one of the main challenges in integrating mobile robots into everyday environments. Applications such as food service, warehouse logistics, and equipment transport could see substantial cost and time savings but require robots to navigate dense human environments with static obstacles.
Social robot navigation research focuses on enabling robots to move safely and efficiently around humans. Deep Reinforcement Learning (DRL) has emerged as a promising alternative to traditional learning-free approaches including reactive, decoupled, and coupled planning. Reactive methods model humans as static objects [1], decoupled planning approaches forecast future human trajectories to generate a cost map for navigation [2], and coupled planning approaches model human motion and solve a joint optimization problem [3]. These methods have notable shortcomings. Reactive approaches are shortsighted, resulting in intrusive behaviour [2]. Decoupled approaches neglect the robot’s impact on humans, which can cause the robot to ‘freeze’ when the robot’s plan conflicts with forecasted human trajectories [4]. Lastly, coupled methods falter when the human model is inaccurate [3], and accurately modeling human motion is challenging.
DRL-based social navigation offers a compelling value proposition: learning efficient and safe robot behaviour without relying on an accurate human behaviour model. Most DRL social navigation research centers around simulation due to the dangers posed by randomly initialized DRL agents and the extensive data required for training [5]. However, these simulators contain human models that are inevitably mismatched with the true dynamics of human motion. Thus, agents trained in these simulators are unfit for deployment. For instance, the popular CrowdNav simulator assumes a cooperative and deterministic human policy [6]: Optimal Reciprocal Collision Avoidance (ORCA) [7]. This assumption leads DRL agents to learn overly aggressive behaviours [8]. Additionally, many simulators such as [6] neglect the presence of static obstacles, further exacerbating the sim-to-real gap.
Figure 1: DR-MPC navigating in the real world. In this illustration, the robot deviates from its path to allow human 1 to pass and then slows down (red means slower speed) to let human 2 to pass before returning to its path.
Figure 2: Full real-world pipeline. The Ouster OS0-128 LiDAR generates a detailed reflectivity image and a point cloud. The reflectivity image allows us to perform human tracking and the point cloud enables localization, path tracking, and depth recovery. With the state constructed from a single sensor, we use our OOD module and CVMM safety check to determine whether or not to execute the DR-MPC or the heuristic safety policy.
In this paper, we train a DRL-based social navigation model directly in the real world. To handle environments with static obstacles, we modify the typical DRL social navigation Markov Decision Process (MDP) into human avoidance with path tracking within virtual corridors. We introduce a novel approach, Deep Residual Model Predictive Control (DR-MPC), that integrates Model Predictive Control (MPC) path tracking to significantly accelerate the learning process. Lastly, we design a pipeline with out-of-distribution (OOD) state detection and a heuristic policy to guide the DRL agent into higher-reward regions. Our approach enabled real-world deployment of a DRL agent without any simulation with less than 4 hours of data.
We begin by discussing related work. Then, we define our decision process formulation in Section III. Next, in Section IV, we detail and motivate our design of DR-MPC. Then, in Section V, we explain how DR-MPC fits into the overall pipeline. In Section VI, we quantitatively demonstrate the effectiveness of our approach in simulation compared to other models. Finally, in Section VII, we present an analysis of our algorithm’s performance when trained on a real robot.
II Related Works
II-A Social Navigation Simulators
Several 2D and 3D simulators have been developed and continue to be developed for social navigation. These simulators offer a way to safely test and experiment with different approaches. Simulators also offer the possibility, in theory, of developing ‘sim-to-real’ approaches, which can substantially reduce the amount of real-world learning (and mistakes) that robots make. A popular simulator for social navigation is the 2D CrowdNav simulator [6], designed for waypoint navigation in open spaces with the human policy defined by ORCA [7]. While CrowdNav has facilitated substantial DRL model development [9, 8, 10, 11], the cooperative nature of the ORCA policy creates a large sim-to-real gap. When the robot is visible to humans, the learned DRL policy is dangerously aggressive: the robot pushes humans out of its way to reach its goal [8]. Consequently, the invisible testbed—when the robot is invisible to the humans—was adopted as the benchmark standard. Although the invisible testbed reduces the DRL agent’s aggression, the problem becomes equivalent to decoupled planning where the DRL agent does not learn how its action will influence humans [8].
Looking beyond CrowdNav, other simulators use alternative human motion models, such as variations of the Social Forces Model (SFM) or behaviour graphs [12, 13, 14]. Unfortunately, even the latest simulators struggle with human realism and often exhibit unsmooth human motions and enter deadlock scenarios. These deficiencies create a significant sim-to-real gap, where models trained in simulation often perform differently–—and usually worse–—when applied in the real world [15]. While these simulators have been essential tools for advancing DRL model architectures, the most accurate data for DRL social navigation is real-world data. By reducing the amount of training data required, we enable the direct training of DR-MPC in the real world, avoiding unnecessary inductive biases introduced by simulators.
Figure 3: DR-MPC architecture. The dark blue text elements involve learning. From $\mathcal{S}_{\text{PT}}$, we generate the MPC path tracking action and a latent embedding of the path information using an MLP. From $\mathcal{S}_{\text{HA}}$, we use a SOTA human avoidance network to generate six actions for human avoidance. The model then fuses all the information to generate $\boldsymbol{\alpha}$ and $\mathbf{p}$, which generates the final action to maximize the human avoidance and path tracking rewards.
II-B RL Social Navigation Models
Crowd navigation, a major sub-field of DRL social navigation, focuses on enabling a robot to navigate among humans in open areas. Significant progress has been made in incorporating machine learning advancements to enhance model architectures that enable reasoning about crowd dynamics in both spatial and temporal dimensions.
Early models analyzed the scene by considering humans individually [16, 17]. Then, approaches incorporated attention mechanisms to reason about the crowd as a whole, using learned attention scores to generate a crowd embedding for decision making [6, 18, 19]. Subsequent advances analyzed both human-robot and human-human interactions through graph neural networks (GNNs) [9, 19, 10]. Most recently, models include temporal reasoning using spatio-temporal graphs, multihead attention mechanisms, and transformers [8, 11], enabling reasoning on the trajectory level.
This paper does not focus on developing a novel architecture for traditional waypoint-based crowd navigation. Instead, we leverage state-of-the-art (SOTA) architectures for processing human-related inputs and incorporate them into DR-MPC.
II-C Residual DRL
Residual DRL integrates two policies together: one is user-supplied, another is learned [20]. DRL learns corrective actions on top of this base controller. The learning speedup occurs because on initialization, the DRL agent will have an expected action equal to the base controller, which, assuming a suitable user-supplied controller, can result in much better performance than a random policy [20].
To the best of our knowledge, residual DRL has not been applied to social navigation, but some related works have attempted to incorporate classical control. For example, Kästner et al. [21] learn a DRL policy that switches between different controllers, each optimized for an individual task. Another work, Semnani et al. [22] switch from a DRL policy to a force-based model when close to humans. While these approaches reduce the learning burden on DRL, they limit the overall performance by the capabilities of the base controllers. When the overall policy selects a base controller, DRL can not optimize the performance of these individual controllers.
In contrast, DR-MPC fully exploits the capabilities of MPC path tracking. DR-MPC can either replicate the MPC action or generate an action that diverges significantly from it. Unlike residual DRL, which follows the base controller actions in expectation, DR-MPC’s initial behaviour almost exactly follows MPC path tracking, representing the best achievable performance without any prior information of humans. Additionally, DR-MPC dynamically learns when to integrate or disregard the MPC action, leading to significantly faster training compared to traditional residual DRL.
II-D Path Tracking
DRL is beneficial for path tracking when modeling difficulties arise, whether related to the agent or environmental disturbances. However, in indoor environments with minimal disturbances and a typical robot that can be accurately modeled with unicycle kinematics, we can leverage MPC for path tracking using the formulation presented by Sehn et al. [23]. As a result, DRL does not need to explicitly learn path tracking, thereby reducing the overall learning complexity.
III Decision process formulation
We modify the social navigation MDP in [6] to combine human avoidance and path tracking within virtual corridors, where these corridors ensure safety from static obstacles. To simplify the problem, we assume a constant corridor width.
III-1 State Space
We construct our state $\mathcal{S}=\{\mathcal{S}_{\text{PT}},\mathcal{S}_{\text{HA}}\}$, where $\mathcal{S}_{\text{PT}}$ is the state information for path tracking and $\mathcal{S}_{\text{HA}}$ is the state information for human avoidance.
As in [8], we exclude human velocities due to the difficulty of estimating these quantities in the real world and assume a constant human radius. If at time $t$ there are $n_{t}$ visible humans: $\mathcal{S}_{\text{HA}}=\left\{\mathbf{v}^{t-H:t-1},\mathbf{r}^{t-H:t},\mathbf{q}^{t-H_{1}:t}_{1},\dots,\mathbf{q}^{t-H_{n_{t}}:t}_{n_{t}}\right\}$, where $\mathbf{v}^{t-H:t-1}$ is the robot’s past velocities from time $t-H$ to $t-1$, $\mathbf{r}^{t-H:t}$ is the robot’s past positions in the current robot frame from time $t-H$ to $t$, and $\mathbf{q}^{t-H_{i}:t}_{i}=\begin{bmatrix}\mathbf{p}^{t-H_{i}}_{i}\dots\mathbf{p}^{t}_{i}\end{bmatrix}$ is the positional trajectory of the $i^{\text{th}}$ visible human in the current robot frame with a look-back of $H_{i}$, capped at length $H$.
We select a local path representation as in [24] to generalize to any path length. Given that $\mathbf{p}_{\ell}$ is the closest localized node to the robot, $\mathcal{S}_{\text{PT}}=[\mathbf{p}_{\ell-L}\ldots\mathbf{p}_{\ell}\ldots\mathbf{p}_{\ell+F}]$, where $L$ is the look-back and $F$ is the look-forward parameter. These path nodes are also transformed into the current robot frame.
III-2 Actions
Our DRL agent interacts at the level of velocity. Our action space is a linear and angular velocity: $\mathbf{a}=\begin{bmatrix}v\hskip 5.69046pt\omega\end{bmatrix}$.
III-3 Rewards
Our reward function considers path advancement, path deviation, goal reaching, corridor collisions, small speeds, human collisions, and human disturbance:
$r=r_{\text{pa}}+r_{\text{dev}}+r_{\text{goal}}^{*}+r_{\text{cor-col}}^{*}+r_{\text{act}}^{*}+r_{\text{hum-col}}^{*}+r_{\text{dist}},$ (1)
where rewards marked with an asterisk are terminal rewards; on all termination conditions, the robot stops moving. All the following $\kappa$ are tuning parameters where $\kappa>0$. Path advancement: $r_{\text{pa}}=\kappa_{\text{pa}}(s^{\prime}_{i}-s_{i})$ where $s_{i}$ and $s^{\prime}_{i}$ are the arclength positions along the path at $\mathcal{S}$ and $\mathcal{S}^{\prime}$. Deviation: $r_{\text{dev}}=\kappa^{\text{xy}}_{\text{dev}}d_{\text{xy}}+\kappa^{\theta}_{\text{dev}}|d_{\theta}|$ where $d_{\text{xy}}$ is the Euclidean distance and $d_{\theta}$ is the angular offset from the closest point on the path. Goal: $r_{\text{goal}}^{*}=-\kappa_{\text{goal}}$ if the heading difference exceeds a threshold and 0 otherwise. Corridor collision: $r_{\text{cor-col}}^{*}=-\kappa_{\text{cor-col}}$ for colliding with the corridor. Minimal actuation: $r_{\text{act}}^{*}=-\kappa_{\text{act}}$ if the sum of the robot’s past $H$ speeds falls below a threshold.
For human avoidance, we align our rewards with the two most important principles of human-robot interaction (HRI): safety and comfort [25]. For safety, human collision: $r_{\text{hum-col}}^{*}=-\kappa_{\text{hum-col}}.$ For comfort, disturbance penalty: $r_{\text{dist}}=-\sum_{i=1}^{n_{t}}\left(\kappa_{\text{v}}|\Delta v_{\text{h}}^{i}|+\kappa_{\theta}|\Delta\theta_{\text{h}}^{i}|\right)$, which penalizes the robot for causing changes in a human’s velocity ($\Delta v_{\text{h}}^{i}$) and heading direction ($\Delta\theta_{\text{h}}^{i}$). These changes are computed relative to the human’s velocity and heading at the previous time step.
Finally, we add two safety layers: safety-human and safety-corridor raises, which are conservative versions of the human-collision and corridor-collision penalties. While a safety violation does not guarantee a collision, it is likely. So, we slightly reduce the theoretical performance limit for safety.
IV DR-MPC Policy Architecture
Our DR-MPC policy architecture (Figure 3) consists of two main components: individual processing of $\mathcal{S}_{\text{PT}}$ and $\mathcal{S}_{\text{HA}}$, and the information fusion to generate a single action.
For processing $\mathcal{S}_{\text{PT}}$, we employ a Multi-layer Perceptron (MLP) to generate a latent embedding of the path information. Also, an MPC optimization guided by [23] generates $\mathbf{a}_{\text{MPC}}$.
To process $\mathcal{S}_{\text{HA}}$, we modify [10] to handle varying-length human trajectories for off-policy learning. We refer to this adapted architecture as the Human Avoidance Network (HAN); further details are provided in the appendix. HAN is a modular component of our pipeline and can be replaced with other crowd navigation work.
The output of HAN is a ‘crowd embedding’ that is used to generate six candidate human avoidance actions ($\mathbf{a}_{\text{HA}}$), where each $\mathbf{a}^{\text{i}}_{\text{HA}}$ is positioned in a different cell of the action space. The human-avoidance action space is partitioned into a grid of six cells defined by two linear velocity bins $[v_{\text{min}},v_{\text{middle}}]$, $[v_{\text{middle}},v_{\text{max}}]$, and three angular velocity bins $[w_{\text{min}},w_{\text{lower}}]$, $[\omega_{\text{lower}},\omega_{\text{upper}}]$, $[\omega_{\text{upper}},\omega_{\text{max}}]$. The inclusion of multiple actions is motivated by the fact that, in many cases, several viable actions may exist for human avoidance. We found empirically that this design reduces learning time compared to having the model learn the diversity. The output of the MLP following HAN is the mean action within each cell. Using a predetermined variance that decays over time, we sample a Gaussian to get $\mathbf{a}^{\text{i}}_{\text{HA}}$; this standard formulation comes from [26].
The key innovation of our model lies in how we combine these individual components. The path tracking embedding, the crowd embedding, $\mathbf{a}_{\text{MPC}}$, and $\mathbf{a}_{\text{HA}}$ are combined to output $\boldsymbol{\alpha}_{\text{raw}}=\begin{bmatrix}\alpha^{1}_{\text{raw}}\dots\alpha^{6}_{\text{raw}}\end{bmatrix}$, where each $\alpha^{i}_{\text{raw}}\in[0,1]$, and a categorical distribution $\mathbf{p}=\begin{bmatrix}p^{1}\dots p^{6}\end{bmatrix}$. Note, as in Soft Actor-Critic (SAC), the model learns the mean and log-standard deviation to generate $\alpha^{i}_{\text{raw}}$, which is then passed through a $\tanh$ function to squash it, followed by scaling and shifting [27]. We then compute $\alpha^{i}=\text{clip}(\alpha^{i}_{\text{raw}}+\tanh(\beta),0,1)$, where $\beta$ is a learned parameter. Each $\mathbf{a}^{\text{i}}_{\text{HA}}$ corresponds to $\alpha^{i}$ and $p^{i}$ of the same index. After sampling the index $j$ from $\mathbf{p}$, the final action is constructed as a weighted sum: $\mathbf{a}=\alpha^{j}\mathbf{a}^{\text{j}}_{\text{HA}}+(1-\alpha^{j})\mathbf{a}_{\text{MPC}}$.
Unlike residual DRL, which follows the base controller in expectation, DR-MPC begins with near-MPC behaviour by biasing actions toward MPC, initializing $\beta=-0.8$ to suppress $\boldsymbol{\alpha}$. Note that initializing $\beta$ too low yields little qualitative difference and requires a lot of model updates to raise $\beta$ towards 0. We observe the model naturally learns to adjust $\beta$ towards 0 to be able to take non-MPC actions.
Although DR-MPC’s effective action space after fusion is smaller than regular DRL, it contains the best actions for path tracking and human avoidance. While DRL with the full action space might theoretically achieve higher performance given infinite training, it faces practical challenges in balancing exploration and exploitation. Despite having a reduced action space, we can still achieve a high-performing policy but with significantly faster learning times.
The DRL algorithm we employ to train our model is Discrete Soft Actor-Critic (DSAC) with double average clipped Q-learning with Q-clip to properly backpropagate through $\mathbf{p}$ [28]. We further augment DSAC with the entropy formulation from SAC to optimize the log-standard deviation that generates $\alpha^{i}_{\text{raw}}$ [27]. Also note that Figure 3 represents the actor model within our actor-critic framework. The critic model shares the same architecture up to and including the first MLP in the fusion stage; however, this MLP outputs the Q-value. The action is included as an input to this MLP for the critic.
V The Pipeline
The exploration-exploitation problem in DRL does not have to be addressed solely by the actor network. To perform better than random exploration, we guide the selection of $\mathbf{a}^{i}_{HA}$ using heuristics to find promising directions for human avoidance. The sooner these valuable experiences are stored in the replay buffer, the sooner the critic can correct these Q-values. This concept of early intervention in uncertain or high-risk states has proven effective in interactive imitation learning and sample-efficient DRL [29, 30].
To implement this, we perform OOD state detection to assess the DRL model’s familiarity with a state. If the state is in-distribution, we execute DR-MPC’s action. Otherwise, we either validate the model’s action as safe using a heuristic safety check or override it using a heuristic policy. This process steers the DRL agent toward collision-free areas.
Once a state is in-distribution, the DRL agent will continue to explore and may still encounter collisions. The agent will quickly learn to focus on known good actions and avoid poor reward regions, thereby accelerating learning speed. Note that the OOD state detection, heuristic safety check, and heuristic policy are modular components and can be swapped out or fine-tuned independently of everything else in the pipeline.
V-1 OOD State Detection
There are numerous methods for OOD state detection: we select a $K$-nearest neighbours (KNN) algorithm in the latent space of the model [31]. This method is relatively interpretable and can be implemented in real-time using Facebook’s Faiss package [32]. We extract the latent embedding from our model at the layer immediately following the fusion of path tracking and human avoidance information.
Figure 4: Simulation results averaged over 12 trials. Both DR-MPC models outperform naïve DRL and Residual DRL. While DRL and Residual DRL learn to achieve a high path advancement reward, they struggle with task switching to perform proficient human avoidance.
V-2 Heuristic Safety Check and Policy
Figure 5: Simulation training scenarios. The path is black, the virutal corridor is red, the robot is yellow, and the humans are light grey.
Upon detecting an OOD state, we assess whether the DRL model’s proposed action is safe. We assume a constant velocity motion model (CVMM) for both the human and the robot. We roll out the proposed action for the robot and the estimated human velocity based on the last two timesteps for a predefined number of timesteps. We then evaluate if the safety-human raise is triggered at any point during this rollout; if it is, the action is deemed unsafe. For the safety-corridor condition, we only check the one-timestep rollout result.
If the proposed DRL action is both OOD and fails the CVMM safety check, we evaluate a predefined set of alternative actions for safety. We select the safe action that is closest to the MPC action, thereby guiding our DRL agent into regions that are both safe and conducive to path advancement.
V-A Soft Resets
In real-world applications, episode resets are inefficient. Thus, unlike traditional episodic RL, we avoid ‘hard resets’ and perform ‘soft resets’ during training by returning the robot to the closest non-terminal state when a termination condition occurs. This approach of starting at any non-terminal state is analogous to the Monte Carlo Exploring Starts algorithm [33].
VI Simulation Experiments
VI-A Simulator and Setup
We augment the CrowdNav simulator from [10], originally developed by [6]. In a modularized manner, we replace waypoint navigation with path tracking. Our training schema cycles through four scenarios, as illustrated in Figure 5. Once a path is completed, the next one begins. We run our simulator with $n_{t}=6$; these 6 humans are modeled using ORCA and continuously move to random goals in the arena. Our robot has an action space of $v\in[0,1]$ and $\omega\in[-1,1]$. The MDP has a timestep of 0.25s, and we train our models with a limited amount of data: 37,500 steps—around 2.5 hours of data.
VI-B Model Baselines
We evaluate four models. The first applies the model of [10], which shares a similar backbone to DR-MPC. The DRL model combines the ‘PT Embedding’ and ‘HA Embedding’ with an MLP to generate the agent’s action; we refer to this model as the naïve DRL model. The second model is a residual DRL model guided by [20], which outputs a corrective action: the action executed in the environment is the sum of the MPC and DRL corrective action, which is then truncated into the feasible action space. Since path tracking and human avoidance may require distinct actions, we set the DRL corrective action to have $v\in[-1,1]$ and $\omega\in[-2,2]$, ensuring that the final action can cover the entire action space. The residual DRL model architecture is the same as the naïve DRL model but also has $\mathbf{a}_{\text{MPC}}$ inputted into the final MLP. Finally, we evaluate two versions of DR-MPC: one without the OOD state detection and CVMM modules, and one with them.
VI-C Results and Discussion
Figure 6: Real-world demonstrations. Yellow is the reference path, orange is the robot’s trajectory, and purple is the human’s trajectory. Left and middle: the robot learns to navigate around humans standing in the robot’s path. Right: The robot learns to deviate from its path to avoid disturbing the human’s trajectory.
The main results are depicted in Figure 4. The plotted lines are the mean values across 12 trials, while the shaded areas indicate the variance observed during training. For all metrics, the values in these plots are averaged over 500 environment steps. We do not use the reward-per-episode metric because different path lengths inherently yield different maximum cumulative rewards, which would skew comparisons.
From the cumulative reward plot, it is clear that both of our DR-MPC models significantly outperform the naïve DRL and residual DRL models. While all models can learn a high path-advancement reward and navigate paths effectively, interacting with humans remains a greater challenge. DR-MPC, however, excels in task switching and optimizing both path tracking and human avoidance using the $\boldsymbol{\alpha}$ parameter. It is important to note that the DRL and residual DRL models were verified for correctness and, given sufficient training, could eventually reach similar reward levels as DR-MPC.
As designed, both DR-MPC models start with a high base reward due to their initialization with near-MPC path tracking behaviour, which is better than the residual DRL model, which follows the MPC action only in expectation, resulting in slower convergence and lower initial rewards.
Comparing our DR-MPC models, the one with OOD state detection performs better in the early stages of training. This advantage comes from the heuristic policy exploring actions deemed safe by the CVMM safety check and that are close to the MPC action, which results in higher path-advancement rewards. The similarity in safety interventions between the two models suggests that ORCA, the human motion model, is not well captured by a CVMM in close contact scenarios. However, the CVMM safety check is valuable in real-world situations, particularly with inattentive humans blindly walking on a straight path or humans who are standing still.
Overall, the results demonstrate that the DR-MPC approach seamlessly integrates MPC behaviour and adjusts to dynamic environments making DR-MPC a robust and efficient solution.
VII Hardware Experiments
VII-A Implementation
VII-A1 Robot
We use a Clearpath Robotics Jackal equipped with an Ouster OS0-128 LiDAR. The OS0 sensor has a $90^{\circ}$ vertical field-of-view (FOV) and a $360^{\circ}$ horizontal FOV, capturing range and reflectivity data. This data can be represented as equirectangular images—an example of the reflectivity image is shown in Figure 2. These images are comparable to low-resolution cameras, enabling the application of computer-vision models. By fusing the 2D computer-vision results with the LiDAR’s depth information, we can produce 3D outputs.
VII-A2 Mapping, Localization, Path Tracking
We use the open-source Teach and Repeat (T&R) codebase [34]. Specifically, in LiDAR T&R, after manually driving the robot through a new environment once, the robot can subsequently localize itself to this previously driven path and track it using MPC. T&R enables rapid deployment of our pipeline in new environments.
Figure 7: DR-MPC’s real-world training results averaged over 6 trials—3 per testing scenario. DR-MPC starts with performance equal to the heuristic policy. However, because of DRL’s ability to learn through noise, after exploration, the final model performs better on each key metric compared to where it started.
Figure 8: Real-world testing scenarios used to evaluate models. These loops contain diverse situations for social navigation.
VII-A3 Human Tracking
We detect humans by running a pre-trained YOLOX model from mmDetection on the reflectivity image and its $180^{\circ}$ shifted version to account for humans at the wrap-around point [35]. By combining depth information from the range image with the localization result from T&R, we recover the 3D world positions of humans in the scene and track them using the Norfair package [36].
VII-B Experimental Setup
During training, between zero and four humans interact with the robot. The MDP timestep is 0.2s, aligned with incoming LiDAR data every 0.1s. Inference is performed on a ThinkPad P16 Gen 1 with an Intel i7-12800HX Processor and NVIDIA RTX A4500 GPU, while real-time training occurs on another computer. Inference model weights are periodically updated.
Although we trained on 3.75 hours of experience transitions, the entire process took about 15 hours. After every 500 experience tuples, we paused data collection to allow the model to update, performing twice as many training updates as environment samples. While off-policy learning can leverage older data, keeping the inference model synchronized with the training model allows Q-values to be corrected as state-action distributions shift. Additionally, outliers caused by processing delays from resource allocation variance were filtered to reduce data noise. Also, soft resets, along with charging the laptop and robot batteries, contributed to the additional time.
We periodically evaluate the model on fixed but diverse scenarios (Figure 8). The scenario on the left involves a stationary human, a human walking toward the robot, free driving, and a crowd crossing. The loop on the right tests the robot’s weaving ability between two stationary humans, followed by free driving, a human crossing the robot’s path, and a differently configured crowd crossing. Each model undergoes 6 runs—3 per scenario.
We benchmark real-world DR-MPC against three models: (1) DR-MPC trained in simulation without modification (‘Sim Model Raw’), (2) DR-MPC trained in simulation with real-world adjustments, such as observation delays, acceleration constraints, and distance-based detection limits (‘Sim Model Adjusted’), and (3) a heuristic policy, which executes $\mathbf{a}_{\text{MPC}}$ if it passes the CVMM safety check and switches to the heuristic policy used to guide the DR-MPC model if unsafe.
The real world introduces challenges absent in simulation, such as perception errors (missed detections, false positives, misidentifications) and variable timing delays from tasks (localization, human detection, action generation). Also, human movements are probabilistic, and the robot faces acceleration constraints based on battery charge. These challenges highlight a key advantage of DRL: it maximizes expected cumulative reward, allowing effective learning in noisy environments.
VII-C Results and Discussion
TABLE I: Real-world Comparison of Final Policies ($\mu\pm\sigma$ Per Loop)
Algorithm Steps to Goal ($\downarrow$) Safety-human Safety-corridor
Raises ($\downarrow$) Raises ($\downarrow$)
Heuristic 158.0 ± 2.7 1.0 ± 0.6 0.3 ± 0.5
Sim Model Raw 175.3 ± 3.4 0.5 ± 0.5 3.2 ± 1.1
Sim Model Adjusted 161.2 ± 4.7 0.7 ± 0.5 1.5 ± 0.8
DR-MPC (Ours) 146.3 ± 3.1 0.3 ± 0.5 0.0 ± 0.0
Table I shows that the final DR-MPC model outperforms benchmark models in all key metrics—fewer safety raises and faster navigation. The training process (Figure 7) begins with performance nearly identical to the heuristic policy (as expected). During exploration, DR-MPC performs worse but eventually coalesces its experience to achieve superior results. After 4 hours of data, DR-MPC’s performance begins to plateau. While more data would add marginal improvements, it is unlikely to completely eliminate safety-human raises. Social navigation is a long-tail problem with rare events, and even in simulation with a deterministic agent, safety-human raises never reach zero.
During training, $\beta$ (orange) trends toward 0, while the average $\alpha$ (blue) remains low, indicating the model learns when to best select the MPC action. There is also an upward trend for in-distribution (green) states. The spike at 5000 steps is not an artifact—early training has major changes in latent representations, making rare state identification difficult, but it quickly resolves with a few thousand environment steps.
Qualitatively, DR-MPC exhibits key crowd-navigation behaviors: smoothly weaving around still humans (left and middle of Figure 6) and deviating from its path and slowing down for moving humans (Figure 1 and right of Figure 6). In human-free areas, DR-MPC switches to MPC path tracking.
The heuristic policy performance suffers because human velocity estimates are noisy, as the position is inferred from the bounding box center, which shifts based on body orientation. Also, this policy does not account for delays or the robot’s acceleration, reducing its ability to navigate around humans. However, it still guides DR-MPC towards open spaces.
‘Sim Model Raw’ performs poorly, oversteering and colliding with virtual boundaries due to unaccounted state delays. Adjusting the simulator (‘Sim Model Adjusted’) improves performance but still falls short of DR-MPC trained in the real world. This underscores the advantage of real-world training in handling noisy data and developing robust policies.
VIII Conclusion
We introduced DR-MPC, a novel integration of MPC path tracking with DRL, and demonstrated its effectiveness and superiority to prior work in both simulation and real-world scenarios. Training a DRL agent directly in the real world bypasses the sim-to-real gap, addressing the inevitable mismatch between the dynamics of modeled humans and real humans. While simulation is crucial for model development, the ultimate goal is deploying DRL agents that perform effectively in real-world conditions, where a plethora of challenges remain.
Acknowledgment
James would like to thank Alexander Krawciw for his invaluable mentorship in working with real robots. James is supported by the Vector Scholarship in AI and the Queen Elizabeth II Graduate Scholarship.
Appendix A Human Avoidance Network
Compared to [10], we use the humans’ past trajectories rather than their forecasted trajectories. This way, DRL directly learns from the sensor noise embedded in the state.
For each human trajectory $\mathbf{q}^{t-H_{i}:t}_{i}=\begin{bmatrix}\mathbf{p}^{t-H_{i}}_{i}\dots\mathbf{p}^{t}_{i}\end{bmatrix}$, we sequentially process it from time $t-H_{i}$ to $t$ using a GRU to generate an embedded trajectory $\mathbf{e}^{i}_{\text{traj}}$. This step handles human trajectories of varying lengths, embedding them into a uniform latent space. Next, with $\mathbf{E}_{\text{traj}}=\begin{bmatrix}\mathbf{e}^{1}_{\text{traj}}\dots\mathbf{e}^{n_{t}}_{\text{traj}}\end{bmatrix}$, we use three MLPs to generate the queries $\mathbf{Q}_{\text{traj}}$, keys $\mathbf{K}_{\text{traj}}$, and values $\mathbf{V}_{\text{traj}}$. We then apply multi-head attention using the scaled dot-product attention: $\text{MultiHead}(\mathbf{Q}_{\text{traj}},\mathbf{K}_{\text{traj}},\mathbf{V}_{\text{traj}})$ (see [37]). The output of this module is $\mathbf{E}_{\text{HH}}$, a $n_{t}\times d_{HH}$ tensor, where $d_{HH}$ is the dimension of the human-human embeddings.
Next, we process the robot trajectory $\mathbf{r}^{t-H:t}$ by embedding it with an MLP into $\mathbf{e}^{\text{robot}}_{\text{traj}}$. We then compute the robot-human attention. Here, the keys $\mathbf{K}_{robot}$ are generated from $\mathbf{e}^{\text{robot}}_{\text{traj}}$ and the queries $\mathbf{Q}_{\text{HH}}$ and the values $\mathbf{V}_{\text{HH}}$ from $\mathbf{E}_{\text{HH}}$. The result of this multi-head attention network is the embedding $\mathbf{e}_{RH}$.
Finally, we concatenate $\mathbf{e}_{RH}$ with $\mathbf{v}^{t-H:t-1}$ and pass this tensor through one last MLP to obtain the crowd embedding $\mathbf{e}_{HA}$, which is then used to generate the 6 mean actions for human avoidance.